The Alford Group had 150,000 shares of common stock outstanding at January 1, 2018. The following activities affected common shares during the year. There are no potential common shares outstanding. 2018 Feb. 28 Purchased 3,000 shares of treasury stock. Oct. 31 Sold the treasury shares purchased on February 28. Nov. 30 Issued 12,000 new shares. Dec. 31 Net income for 2018 is $1,639,000. 2019 Jan. 15 Declared and issued a 2-for-1 stock split. Dec. 31 Net income for 2019 is $1,639,000. Required: 1. Determine the 2018 EPS. 2. Determine the 2019 EPS. 3. At what amount will the 2018 EPS be presented in the 2019 comparative financial statements? (For all requirements, Enter your answers in thousands.)
Answer #1

1)

Period Transaction Number of shares outstanding period of outstanding out of 12 months weighted average number of shares outstanding
1Jan 2018 Beginning 150000 2/12 [1JAN-28FEB] 150000*2/12=25000
Feb 28 Purchase of treasury stock 150000-3000=147000 8/12   [1MAR-31OCT] 147000*8/12= 98000
Oct 31 Reissue of shares 147000+3000=150000 1/12    [1NOV-30NOV] 150000*1/12=12500
Nov 30 issue of shares 150000+12000=162000 1/12   [1DEC-31DEC] 162000*1/12 = 13500
Weighted average number of shares outstanding 149000

Earning per share = [Net income -preferred dividend if any ]/Weighted average number of shares outstanding

                 = [1639000 -0]/149000

                = $ 11 per share

2)In case of stock split ,It is assumed to be made at beginning of period .

Number of shares outstanding throughout 2019 = 162000 *2 /1= 324000 shares

Earning per share = [Net income -preferred dividend if any ]/Weighted average number of shares outstanding

                 = [1639000 -0]/324000

                = $ 5.06 per share

3)

For comparative purpose ,It is assumed that Stock split is made at beginning of 2018 .so weighted average number of shares outstanding in 2018 = 149000 (as calculated in part a) *2 /1 = 298000 shares

Earning per share for 2018= [Net income -preferred dividend if any ]/Weighted average number of shares outstanding

                 = [1639000 -0]/298000

                = $ 5.5 per share

2019 = $ 5.06 per share